OpenStudy (anonymous):

find the missing coordinate......20x + 3y=27 (1/2,y)

OpenStudy (anonymous):

x=1/2, put that value in, see what happens.

OpenStudy (anonymous):

so 20 * 1/2 =10

OpenStudy (anonymous):

right, so we get 10+3y=27

OpenStudy (anonymous):

so 1/2 represents y also

OpenStudy (anonymous):

No, you need to solve 10+3y=27 to get y.

OpenStudy (anonymous):

ok you put a multiple with 3 to get your answer?

OpenStudy (anonymous):

ehh, what? Do you know how to solve 10+3y=27?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Let's do it together.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

first we need to get the 10 on the other side, how do we do that?

OpenStudy (anonymous):

10+3(10)

OpenStudy (anonymous):

I'm afraid not, subtract 10 from both sides of the equation.

OpenStudy (anonymous):

(10-10)+ (3-10)

OpenStudy (anonymous):

10 - 3y is one side of the equation, 27 is the other. So when we substract 10 from both sides we get 10-3y-10=27-10 Right?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what does 10-3y-10 equal?

OpenStudy (anonymous):

?????

OpenStudy (anonymous):

10-3y-10=10-10-3y=?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

10-3y-10= -3y and 10-10-3y=-3y...... so what am I doing wrong

OpenStudy (anonymous):

right -3y

OpenStudy (anonymous):

-3y=27-10

OpenStudy (anonymous):

27-10=17 so the equation we get is -3y=17 right?

OpenStudy (anonymous):

;~)

OpenStudy (anonymous):

we're almost there, how do we get to y=... ?

OpenStudy (anonymous):

wow i'm going to fail my test I thought we were done

OpenStudy (anonymous):

You're done when you get to y=... now we have -3y=17

OpenStudy (anonymous):

You need to divide both sides of the equation by -3 to get to y=...

OpenStudy (anonymous):

what does the left side of the equation become if we divide by -3?

OpenStudy (anonymous):

if i'm dividing -3 into -3????

OpenStudy (anonymous):

you divide -3y by -3

OpenStudy (anonymous):

-y

OpenStudy (anonymous):

what's \[\frac{-1}{-1}\]?

OpenStudy (anonymous):

i can't figuire it out

OpenStudy (anonymous):

\[\frac{-1}{-1}=1\] if you divide a negetive number by a negative number you get a positive number.

OpenStudy (anonymous):

so what's \[\frac{-3y}{-3}\]

OpenStudy (anonymous):

6y

OpenStudy (anonymous):

\[\frac{2}{2}=1, \frac{-1}{-1}=1, \frac{-3}{-3}=1\]

OpenStudy (anonymous):

\[\frac{-3y}{-3}=\frac{-3}{-3}y\]

OpenStudy (anonymous):

So \[\frac{-3y}{-3}=y\] right?

OpenStudy (anonymous):

??????

OpenStudy (anonymous):

what is 2y/2?

OpenStudy (anonymous):

are they both positive

OpenStudy (anonymous):

yes, there are no minus signs.

OpenStudy (anonymous):

y is the answer

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so what's -3y/-3

OpenStudy (anonymous):

y

OpenStudy (anonymous):

that's right

OpenStudy (anonymous):

that's one side of the equation

OpenStudy (anonymous):

now the other 17/-3

OpenStudy (anonymous):

i have just lost the whole purpose of the equation

OpenStudy (anonymous):

We started out with 20x+3y=27, x was 1/2 and to find y we had to solve the equation 20*1/2 +3y=27

OpenStudy (anonymous):

which makes me see that we made a mistake.

OpenStudy (anonymous):

we had -3y, but it has to be 3y

OpenStudy (anonymous):

can you write it in long form

OpenStudy (anonymous):

You mean \[\frac{a}{b}\] what do you want me to write?

OpenStudy (anonymous):

this is very long I have to put this on the back burning and move to the next problem

OpenStudy (anonymous):

Well y=17/3 for you information.

OpenStudy (anonymous):

thank you for your help

OpenStudy (anonymous):

You're welcome.